OpenStudy (anonymous):

Dakota’s math test grade was 7 points less than his science test grade. The sum of the grades was 183%. What did Dakota get on his math test? Select one: a. 83% b. 88% c. 93% d. 95%

OpenStudy (anonymous):

do 183%/2

OpenStudy (anonymous):

then take that awnser and make it twice for example if it was 23 do 23 and 23 then when u add 1 to one of them subtract the other such as 22 and 24 do that until you get them to be seven away from each other

OpenStudy (anonymous):

then the one with the lowest amount will be the math grade

OpenStudy (misssunshinexxoxo):

Science Test Grade = x. Maths Test Grade = x - 7. 183 = x + x - 7 183 = 2x - 7 190 = 2x 190/2 = x x = 95 (%).